Give examples of three enzyme-catalyzed reactions and discuss each one.
Answer #1
  1. Alcohol dehydrogenase converts primary alcohols to aldehydes.

HCCH OH + NAD HACCHO + NADH+H

In this reaction, ethanol is converted to acetaldehyde, and the cofactor, NAD, is converted to NADH. In other words, ethanol is oxidized, and NAD is reduced. (The charges don't balance, because NAD has some other charged groups.) Remember that in redox reactions, one substrate is oxidized and one is reduced.

2. Alanine aminotransferase shuffles the alpha‐amino group between alanine and aspartate:

COOCOO HN-C-H + O= C CH2 COO COO + HINCH CH3 CH3 CH2 COO coo Alanine Oxaloacetate Pyruvate Aspartate

Other transferases move phosphate groups between ATP and other compounds, sugar residues to form disaccharides, and so on.

3. Phosphatases break the oxygen‐phosphorus bond of phosphate esters:



R-O-P- 0 1 R-OH -0-0 HO-P-o

Other hydrolases function as digestive enzymes, for example, by breaking the peptide bonds in proteins.

4. Decarboxylases remove CO 2 from alpha‐ or beta‐keto acids:

Dehydratases remove water, as in fumarase (fumarate hydratase):


CO2 8-ي=-3 HO-C-H H-C-H CO2

Deaminases remove ammonia, for example, in the removal of amino groups from amino acids:


CO2 CO2 + NHS HEN-C-H CH OH C=O CH3 Serine Pyruvate

I think 4 example should be enough